# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 3.2 Projective spaces over the full matrix ring
Here we consider our first main example – the projective space over the full matrix ring, cf. [24, Section 7] and [22].
Let $\mathbb{F}$ be a field, and $R=\text{M}_{k}(\mathbb{F})$ be the ring of $k\times k$ matrices over $\mathbb{F}$. By definition, projective spaces over $R$ are projectivizations of $R$-modules. Since $R$ is isomorphic to its opposite ring via transposition, there is no distinction between left and right modules. So, we will restrict our attention to left modules. Any left module $M$ over $\text{M}_{k}(\mathbb{F})$ is isomorphic to the one of the form $V^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}V$, where $V$ is a (not necessarily finite-dimensional) vector space over $\mathbb{F}$, and $\text{M}_{k}(\mathbb{F})$ acts on the first tensor factor. The correspondence $V\mapsto V^{k}$ provides an additive equivalence between the categories of $\mathbb{F}$-modules (vector spaces) and $\text{M}_{k}(\mathbb{F})$-modules. This is also known as the Morita equivalence between the corresponding rings $\mathbb{F}$ and $\text{M}_{k}(\mathbb{F})$. Thus, for the purpose of describing projective spaces over $\text{M}_{k}(\mathbb{F})$, it suffices to consider left modules of the form $M=V^{k}$, where $V$ is a vector space over $\mathbb{F}$.
@@
Submodules of the module $V^{k}$ are of the form $W^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}W$ where $W\subset V$ is an arbitrary subspace of $V$. The correspondence $W\mapsto W^{k}$ gives an isomorphism between the lattice of subspaces of $V$ and the lattice of submodules of $V^{k}$. In particular, any submodule of $V^{k}$ admits a direct complement, which means that subspaces of $V^{k}$ are precisely its free submodules. Further, $W^{k}$ is a free rank one $\text{M}_{k}(\mathbb{F})$-module precisely when $\dim_{\mathbb{F}}W=k$, so rank one free submodules of $V^{k}$ are those of the form $W^{k}$ for $W\in\mathrm{Gr}(k,V)$. ∎
@@
By Corollary 3.8, the space $\mathbb{P}(M)$ can alternatively be described as the set of unimodular elements in $M$ modulo units. For a vector space $V$ of finite dimension $n$, the module $V^{k}$ is isomorphic to the module $\text{M}_{k,n}(\mathbb{F})$ of $k\times n$ matrices. Such a matrix is unimodular in the sense of Definition 3.5 if and only if it has full rank. Thus
@@ 3.3 The $d$-dimensional projective space
Suppose that $M$ is (left or right) free $R$-module of rank $d+1$. Then we call $\mathbb{P}(M)$ a (respectively, left or right) projective space of dimension $d$.
We have two kinds of projective spaces: left projective spaces (projectivizations of left modules) and right projective spaces (projectivizations of right modules). In the rest of this section, we restrict our attention to left projective spaces. The case of right projective spaces is considered analogously. Clearly, all (left) projective spaces of the same (finite) dimension are projectively equivalent (i.e., are related to each other by a projective transformation). We will frequently not distinguish between projectively equivalent objects and talk about the $d$-dimensional projective space over $R$, denoted $\mathbb{P}^{d}(R)$. Without loss of generality, one can think of $\mathbb{P}^{d}(R)$ as the projectivization of $R^{d+1}$.
@@
An incidence structure is a set of points, denoted by $\mathcal{P}$, a set of lines, denoted $\mathcal{L}$, and an incidence $\mathcal{I}\subset\mathcal{P}\times\mathcal{L}$. We say $P\in\mathcal{P},L\in\mathcal{L}$ are incident and write $P\in L$ if $(P,L)\in\mathcal{I}$. Two incidence structures $(\mathcal{P},\mathcal{L},\mathcal{I})$ and $(\mathcal{P}^{\prime},\mathcal{L}^{\prime},\mathcal{I}^{\prime})$ are isomorphic if there are bijections $\mathcal{P}\to\mathcal{P}^{\prime}$, $\mathcal{L}\to\mathcal{L}^{\prime}$ such that the image of $\mathcal{I}$ under the induced map $\mathcal{P}\times\mathcal{L}\to\mathcal{P}^{\prime}\times\mathcal{L}^{\prime}$ is precisely $\mathcal{I}^{\prime}$.
If $\mathcal{P}$ is a point set of some (fixed) incidence structure, we will say that $\mathcal{P}$ is an incidence space. We will say that a map $\mathcal{P}\to\mathcal{P}^{\prime}$ is an isomorphism of incidence spaces if it extends to an isomorphism of incidence structures.
@@
$\mathcal{L}(\mathbb{P}(M)):=\{\mbox{rank two subspaces of $M$}\},$
and the incidence $\mathcal{I}\subset\mathbb{P}(M)\times\mathcal{L}(\mathbb{P}(M))$ is given by containment ($(P,L)\in\mathcal{I}\iff P\subset L$).
@@ 3.4 The projective plane over the full matrix ring
Back to the example $R=\text{M}_{k}(\mathbb{F})$, notice that $V^{k}$ is a free $\text{M}_{k}(\mathbb{F})$-module of rank $d+1$ if and only if $\dim V=k(d+1)$. Thus, by Proposition 3.9, we have
$\mathbb{P}^{d}(\text{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,(d+1)k).$
@@
$\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,3k).$
Furthermore, the space of lines in $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows.
@@
There is a natural identification $\mathcal{L}(\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F})))\simeq\mathrm{Gr}(2k,3k)$. The incidence relation between points $P\in\mathrm{Gr}(k,3k)$ and lines $L\in\mathrm{Gr}(2k,3k)$ is given by containment.
@@
We have $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. Recall that the correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with lattice of submodules of $V^{k}$. Furthermore, $W^{k}$ is free of rank two if and only if $\dim W=2k$. Thus, lines in $\mathbb{P}(V^{k})$ correspond to $2k$-dimensional subspaces of $V$, and incidence between points and lines is given by containment. ∎
In words, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows: points are $k$-dimensional subspaces of $\mathbb{F}^{3k}$, lines are $2k$-dimensional subspaces of $\mathbb{F}^{3k}$, while incidence of points and lines corresponds to containment. Note that two generic points $P_{1},P_{2}\in\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ lie on a unique line $P_{1}\oplus P_{2}$. Likewise, two generic lines $L_{1},L_{2}$ meet at a unique point $L_{1}\cap L_{2}$. However, if $\dim P_{1}\oplus P_{2}<2k$, then there exist infinitely many lines containing points $P_{1},P_{2}$ (and, if $\dim L_{1}\cap L_{2}>k$, there exist infinitely many points lying on $L_{1},L_{2}$). Thus, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is not a combinatorial projective plane. Incidence spaces with properties similar to those of $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ are known as Barbilian planes, see [3, 23]. It is shown in [23] that such planes are precisely the projective planes over rings of stable rank two, of which the matrix ring is an example.
@@ 3.5 The projective plane over the dual numbers
Let $\mathbb{D}:=\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ be the ring of dual numbers. In this section, we show that the geometry of $\mathbb{P}^{2}(\mathbb{D})$ is the skewer geometry of [20], where both points and lines are affine lines in $\mathbb{R}^{3}$, and a point is incident to a line if and only if the corresponding affine lines meet at a right angle.
@@
Let $\mathcal{A}(V)$ be the space of affine straight lines in $V$. We define an incidence structure on $\mathcal{A}(V)$ as follows: both points and lines are affine lines in $V$, i.e., $\mathcal{P}=\mathcal{L}=\mathcal{A}(V)$, and a point is incident to a line if and only if the affine lines intersect and are perpendicular to each other. In this geometry, both the operation of joining two points with a line and intersecting two lines boil down to taking the common perpendicular, or skewer, of the corresponding affine lines.
We will show that there is natural isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$. The corresponding mapping $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ between the point spaces is a variant of the Study mapping, see [16, Section 2.3] and [19]. For an affine line $\{p+tv\mid t\in\mathbb{R}\}$, where $p,v\in V$ and $v\neq 0$, let
@@
where the notation $\times$ stands the cross product in $V$, and define
$\phi_{P}(p+tv):=\mathbb{D}\zeta(v,p).$
@@
The mapping $\phi_{P}$ is well-defined and bijective.
@@
In particular, for any $v\neq 0$ we have that $\zeta(v,p)$ is unimodular and therefore $\mathbb{D}\zeta(v,p)$ is a subspace. Moreover, this subspace does not depend on the choice of the direction vector $v$ and base point $p$ of the line $v+tp$, since for any $\lambda\in\mathbb{R}$ we have $\zeta(\lambda v,p)=\lambda\zeta(v,p)$ and $\zeta(v,p+\lambda v)=\zeta(v,p)$. So, the mapping $\phi_{P}$ is well-defined.
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace of rank $k$. Then $N^{\bot}:=\{v\in M\mid\langle v,w\rangle=0\mbox{ for all }w\in N\}$ is also a subspace, and $\mathrm{rank}\,N^{\bot}=n-k$.
@@
Consider the restriction mapping $\rho\colon M^{*}\to N^{*}$. Since $N$ is free of rank $k$, so is $N^{*}$, and $M^{*}=\mathrm{Ker}\,\rho\oplus K$ for some submodule $K\subset M^{*}$. Thus, by Proposition 3.17, the submodule $\mathrm{Ker}\,\rho\subset M{{}^{*}}$ is free. Furthermore, $\rho_{K}\colon K\to N^{*}$ is an isomorphism, so $K$ has rank $k$ and $\mathrm{Ker}\,\rho$ has rank $n-k$. Now, consider the mapping $\phi\colon M\to M^{*}$ given by the inner product. Since the inner product is non-degenerate, the mapping $\phi$ is injective. Therefore, since $M$ is a finite-dimensional real algebra, $\phi$ is, in fact, an isomorphism. Together, $N^{\bot}=\phi^{-1}(\mathrm{Ker}\,\rho)$ is a free submodule of $M$ which has rank $n-k$ and admits a direct complement, $\phi^{-1}(K)$, as needed. ∎
@@
Suppose $M$ is a free $\mathbb{D}$-module of finite rank, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace. Then $(N^{\bot})^{\bot}=N$.
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Then the correspondence $N\mapsto N^{\bot}$ is a bijection:
{subspaces of $\{\text{subspaces of }M\mbox{of rank }k\}\leftrightarrow\{\text{subspaces of }M\mbox{of rank }n-k\}.$
We now define a map $\phi_{L}\colon\mathcal{A}(V)\to\mathcal{L}(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ by
$\phi_{L}(p+tv):=\phi_{P}(p+tv)^{\bot},$
where the inner product on $\mathbb{D}\otimes_{\mathbb{R}}V$ is defined as the unique $\mathbb{D}$-linear extension of the inner product on $V$. It follows from Proposition 3.16 and Corollary 3.20 that the map $\phi_{L}$ is well-defined and bijective.
@@
The mappings $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ and $\phi_{L}\colon\mathcal{A}(V)\to\mathcal{L}(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ provide an isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$.
@@
We already know that the mappings $\phi_{P}$, $\phi_{L}$ are bijections, so it remains to show that they preserve incidence. In other words, we need to show that the lines $p+tv$ and $q+tw$ meet at a a right angle if and only if $\phi_{P}(p+tv)\subset\phi_{L}(q+tw)$. We have
$\phi_{P}(p+tv)\subset\phi_{L}(q+tw)=\phi_{P}(q+tw)^{\bot}\quad\Leftrightarrow\quad\langle\zeta(p+tv),\zeta(q+tw)\rangle=0.$
Further,
@@
But $\det(v,w,q-p)=0$ if either $v$ and $w$ are collinear (which can not happen if $\langle v,w\rangle=0$, since the inner product on $V$ is positive-definite), or $q-p\in\text{span}\langle v,w\rangle$. The latter is equivalent to saying that the lines $p+tv$ and $q+tw$ have a point in common. So, combining this with the condition $\langle v,w\rangle=0$, we get that $\phi_{P}(p+tv)\subset\phi_{L}(q+tw)$ if and only if the lines $p+tv$ and $q+tw$ meet at a right angle, as needed. ∎

### v3

@@ 3.2 Projective spaces over the full matrix ring
Here we consider our first main example – the projective space over the full matrix ring, cf. [29, Section 7] and [26].
Let $\mathbb{F}$ be a field, and $R=\mathrm{M}_{k}(\mathbb{F})$ be the ring of $k\times k$ matrices over $\mathbb{F}$. By definition, projective spaces over $R$ are projectivizations of $R$-modules. Since $R$ is isomorphic to its opposite ring via transposition, there is no distinction between left and right modules. So, we will restrict our attention to left modules. Any left $\mathrm{M}_{k}(\mathbb{F})$-module $M$ is isomorphic to the one of the form $V^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}V$, where $V$ is a (not necessarily finite-dimensional) vector space over $\mathbb{F}$, and $\mathrm{M}_{k}(\mathbb{F})$ acts on the first tensor factor [^1] Note that, as an $\mathbb{F}$-vector space, $V^{k}$ is just a direct sum of $k$ copies of $V$, hence the notation. We define it as $\mathbb{F}^{k}\otimes_{\mathbb{F}}V$ to emphasize the $\mathrm{M}_{k}(\mathbb{F})$-module structure. The correspondence $V\mapsto V^{k}$ provides an additive equivalence between the categories of $\mathbb{F}$-modules (vector spaces) and $\mathrm{M}_{k}(\mathbb{F})$-modules. This is known as the Morita equivalence between the corresponding rings $\mathbb{F}$ and $\mathrm{M}_{k}(\mathbb{F})$.. Thus, for the purpose of describing projective spaces over $\mathrm{M}_{k}(\mathbb{F})$, it suffices to consider left modules of the form $M=V^{k}$, where $V$ is a vector space over $\mathbb{F}$.
@@
Submodules of the module $V^{k}$ are of the form $W^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}W$ where $W\subset V$ is an arbitrary subspace of $V$. The correspondence $W\mapsto W^{k}$ gives an isomorphism between the lattice of subspaces of $V$ and the lattice of submodules of $V^{k}$. In particular, any submodule of $V^{k}$ admits a direct complement, which means that subspaces of $V^{k}$ are precisely its free submodules. Further, $W^{k}$ is a free rank one $\mathrm{M}_{k}(\mathbb{F})$-module precisely when $\dim_{\mathbb{F}}W=k$, so rank one free submodules of $V^{k}$ are those of the form $W^{k}$ for $W\in\mathrm{Gr}(k,V)$. ∎
@@
By Corollary 3.10, the space $\mathbb{P}(M)$ can alternatively be described as the set of unimodular elements in $M$ modulo units. For a vector space $V$ of finite dimension $n$, the module $V^{k}$ is isomorphic to the module $\mathrm{M}_{k,n}(\mathbb{F})$ of $k\times n$ matrices. Such a matrix is unimodular in the sense of Definition 3.6 if and only if it has full rank. Thus
@@ 3.3 The d-dimensional projective space
Suppose that $M$ is left free $R$-module of rank $d+1$. Then we call $\mathbb{P}(M)$ a (left) projective space of dimension $d$.
Clearly, all (left) projective spaces of the same (finite) dimension are projectively equivalent (i.e., are related to each other by a projective transformation). We will frequently not distinguish between projectively equivalent objects and talk about the $d$-dimensional projective space over $R$, denoted $\mathbb{P}^{d}(R)$. Without loss of generality, one can think of $\mathbb{P}^{d}(R)$ as the projectivization of $R^{d+1}$.
@@
An incidence structure is a set of points, denoted by $\Pi$, a set of lines, denoted $\Lambda$, and an incidence $\mathrm{I}\subset\Pi\times\Lambda$. We say $P\in\Pi,L\in\Lambda$ are incident and write $P\in L$ if $(P,L)\in\mathrm{I}$. Two incidence structures $(\Pi,\Lambda,\mathrm{I})$ and $(\Pi^{\prime},\Lambda^{\prime},\mathrm{I}^{\prime})$ are isomorphic if there are bijections $\Pi\to\Pi^{\prime}$, $\Lambda\to\Lambda^{\prime}$ such that the image of $\mathrm{I}$ under the induced map $\Pi\times\Lambda\to\Pi^{\prime}\times\Lambda^{\prime}$ is precisely $\mathrm{I}^{\prime}$.
If $\Pi$ is a point set of some (fixed) incidence structure, we will say that $\Pi$ is an incidence space. We will say that a map $\Pi\to\Pi^{\prime}$ is an isomorphism of incidence spaces if it extends to an isomorphism of incidence structures.
@@
$\Lambda(\mathbb{P}(M)):=\{\mbox{rank two subspaces of $M$}\},$
and the incidence $\mathrm{I}\subset\mathbb{P}(M)\times\Lambda(\mathbb{P}(M))$ is given by containment ($(P,L)\in\mathrm{I}\iff P\subset L$).
@@ 3.4 The projective plane over the full matrix ring
Back to the example $R=\mathrm{M}_{k}(\mathbb{F})$, notice that $V^{k}$ is a free $\mathrm{M}_{k}(\mathbb{F})$-module of rank $d+1$ if and only if $\dim V=k(d+1)$. Thus, by Proposition 3.11, we have
$\mathbb{P}^{d}(\mathrm{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,(d+1)k).$
@@
$\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,3k).$
Furthermore, the space of lines in $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ can be described as follows.
@@
There is a natural identification $\Lambda(\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F})))\simeq\mathrm{Gr}(2k,3k)$. The incidence relation between points $P\in\mathrm{Gr}(k,3k)$ and lines $L\in\mathrm{Gr}(2k,3k)$ is given by containment.
@@
We have $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. Recall that the correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with lattice of submodules of $V^{k}$. Furthermore, $W^{k}$ is free of rank two if and only if $\dim W=2k$. Thus, lines in $\mathbb{P}(V^{k})$ correspond to $2k$-dimensional subspaces of $V$, and incidence between points and lines is given by containment. ∎
In words, $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ can be described as follows: points are $k$-dimensional subspaces of $\mathbb{F}^{3k}$, lines are $2k$-dimensional subspaces of $\mathbb{F}^{3k}$, while incidence of points and lines corresponds to containment. Note that two generic points $P_{1},P_{2}\in\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ lie on a unique line $P_{1}\oplus P_{2}$. Likewise, two generic lines $L_{1},L_{2}$ meet at a unique point $L_{1}\cap L_{2}$. However, if $\dim P_{1}\oplus P_{2}<2k$, then there exist infinitely many lines containing points $P_{1},P_{2}$ (and, if $\dim L_{1}\cap L_{2}>k$, there exist infinitely many points lying on $L_{1},L_{2}$). Thus, $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ is not a combinatorial projective plane. Incidence spaces with properties similar to those of $\mathbb{P}^{2}(\mathrm{M}_{k}(\mathbb{F}))$ are known as Barbilian planes, see [4, 28]. It is shown in [28, Section 5.12] that such planes are precisely the projective planes over rings of stable rank two, of which the matrix ring is an example.
@@ 3.5 The projective plane over the dual numbers
Let $\mathbb{D}:=\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ be the ring of dual numbers. In this section, we show that the geometry of $\mathbb{P}^{2}(\mathbb{D})$ is the skewer geometry of [24], where both points and lines are affine lines in $\mathbb{R}^{3}$, and a point is incident to a line if and only if the corresponding affine lines meet at a right angle. We note that the relationship between dual numbers and lines in $\mathbb{R}^{3}$ has been known since the introduction of the former; see, e.g., [19, Section 2.3] and [24, Section 2.6]. However, quite strikingly, the result on the precise coincidence of the corresponding geometries does not seem to appear in the existing literature.
@@
Let $\mathcal{A}(V)$ be the space of affine straight lines in $V$. We define an incidence structure on $\mathcal{A}(V)$ as follows: both points and lines are affine lines in $V$, i.e., $\Pi=\Lambda=\mathcal{A}(V)$, and a point is incident to a line if and only if the affine lines intersect and are perpendicular to each other. In this geometry, both the operation of joining two points with a line and intersecting two lines boil down to taking the common perpendicular, or skewer, of the corresponding affine lines.
We will show that there is natural isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$. The corresponding mapping $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ between the point spaces is a variant of the Study mapping, see [19, Section 2.3] and [23]. For an affine line $\{p+tv\mid t\in\mathbb{R}\}$, where $p,v\in V$ and $v\neq 0$, let
@@
where the notation $\times$ stands for the cross product in $V$, and define
$\phi_{\mathrm{P}}(p+tv):=\mathbb{D}\zeta(v,p).$
@@
The mapping $\phi_{\mathrm{P}}$ is well-defined and bijective.
@@
In particular, for any $v\neq 0$ we have that $\zeta(v,p)$ is unimodular and therefore $\mathbb{D}\zeta(v,p)$ is a subspace. Moreover, this subspace does not depend on the choice of the direction vector $v$ and base point $p$ of the line $v+tp$, since for any $\lambda\in\mathbb{R}$ we have $\zeta(\lambda v,p)=\lambda\zeta(v,p)$ and $\zeta(v,p+\lambda v)=\zeta(v,p)$. So, the mapping $\phi_{\mathrm{P}}$ is well-defined.
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-bilinear inner product. Let $N\subset M$ be a subspace of rank $k$. Then $N^{\bot}:=\{v\in M\mid\langle v,w\rangle=0\mbox{ for all }w\in N\}$ is also a subspace, and $\mathrm{rank}\,N^{\bot}=n-k$.
@@
Consider the restriction mapping $\rho\colon M^{*}\to N^{*}$. Since $N$ is free of rank $k$, so is $N^{*}$, and $M^{*}=\mathrm{Ker}\,\rho\oplus K$ for some submodule $K\subset M^{*}$. Thus, by Proposition 3.19, the submodule $\mathrm{Ker}\,\rho\subset M{{}^{*}}$ is free. Furthermore, $\rho_{K}\colon K\to N^{*}$ is an isomorphism, so $K$ has rank $k$ and $\mathrm{Ker}\,\rho$ has rank nkn-k. Now, consider the mapping $\phi\colon M\to M^{*}$ given by the inner product. Since the inner product is non-degenerate, the mapping $\phi$ is injective. Therefore, since $M$ is a finite-dimensional real algebra, $\phi$ is, in fact, an isomorphism. Together, $N^{\bot}=\phi^{-1}(\mathrm{Ker}\,\rho)$ is a free submodule of $M$ which has rank nkn-k and admits a direct complement, $\phi^{-1}(K)$, as needed. ∎
@@
Suppose $M$ is a free $\mathbb{D}$-module of finite rank, equipped with a non-degenerate $\mathbb{D}$-bilinear inner product. Let $N\subset M$ be a subspace. Then $(N^{\bot})^{\bot}=N$.
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-bilinear inner product. Then the correspondence $N\mapsto N^{\bot}$ is a bijection:
[…]
We now define a map $\phi_{\mathrm{L}}\colon\mathcal{A}(V)\to\Lambda(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ by
$\phi_{\mathrm{L}}(p+tv):=\phi_{\mathrm{P}}(p+tv)^{\bot},$
where the inner product on $\mathbb{D}\otimes_{\mathbb{R}}V$ is defined as the unique $\mathbb{D}$-bilinear extension of the inner product on $V$. It follows from Proposition 3.18 and Corollary 3.22 that the map $\phi_{\mathrm{L}}$ is well-defined and bijective.
@@
The mappings $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ and $\phi_{\mathrm{L}}\colon\mathcal{A}(V)\to\Lambda(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ provide an isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$.
@@
We already know that the mappings $\phi_{\mathrm{P}}$, $\phi_{\mathrm{L}}$ are bijections, so it remains to show that they preserve incidence. In other words, we need to show that the lines $p+tv$ and $q+tw$ meet at a a right angle if and only if $\phi_{\mathrm{P}}(p+tv)\subset\phi_{\mathrm{L}}(q+tw)$. We have
$\phi_{\mathrm{P}}(p+tv)\subset\phi_{\mathrm{L}}(q+tw)=\phi_{\mathrm{P}}(q+tw)^{\bot}\quad\Leftrightarrow\quad\langle\zeta(p+tv),\zeta(q+tw)\rangle=0.$
Furthermore,
@@
But $\det(v,w,q-p)=0$ if either $v$ and $w$ are collinear (which can not happen if $\langle v,w\rangle=0$, since the inner product on $V$ is positive-definite), or $q-p\in\text{span}\langle v,w\rangle$. The latter is equivalent to saying that the lines $p+tv$ and $q+tw$ have a point in common. So, combining this with the condition $\langle v,w\rangle=0$, we get that $\phi_{\mathrm{P}}(p+tv)\subset\phi_{\mathrm{L}}(q+tw)$ if and only if the lines $p+tv$ and $q+tw$ meet at a right angle, as needed. ∎